# Evaluating Interventional Reasoning Capabilities of Large Language Models

### v1

@@ 4.2 Empirical studies on benchmarks
RQ2: To what extent is LLM performance affected by possibly memorized causal relations?  While we might be tempted to conclude that GPT-4 reliably predicts changes to models after interventions are performed, we consider spurious factors that can affect model performance. In particular, Kıcıman et al. (2023) found that GPT models reliably retrieved information about TP causal relations, suggesting that these relations could have been included in the training data for LLMs. (We reproduce their findings in Appendix Table 6.) This leads to a worrying possibility: after interventions, could LLMs fail to update their beliefs about causal relations that they have potentially memorized?
To study this RQ, we consider the anti-commonsense benchmark, $\mathcal{D}_{3}$. We consider the intervention effects defined based on: 1) for the bivariate DAG in Figure 1, consider an intervention on $B$ and consider the causal relations $C_{AB}$; for the mediation DAG in Figure 1, consider interventions on $B$ and on $C$, and causal relations $C_{AC}$. Crucially, these intervention effects were chosen so that if an LLM attends only to the plausibly memorized causal relation, e.g., $C_{AB}(G)$ in the bivariate DAG $G$, then under intervention to $B$, where the causal relation $C_{AB}(G^{B})$ changes, the LLM will be penalized if it relies on its memorized fact. This reasoning holds true for intervention effects defined the mediation DAG as well.
We contrast the performance that LLMs achieve on this anti-commonsense benchmark against their performance on the Tübingen benchmark, $\mathcal{D}_{2}$. In $\mathcal{D}_{2}$, although the names of variables come from the same distribution (the TP dataset), the target causal relations such as $C_{AB}$ are not ground-truth relations in TP, meaning that LLMs cannot rely on their memory when being evaluated on this benchmark. If we saw no statistically significant differences in prediction when contrasting the same exact intervention effect but with prompts instantiated according $\mathcal{D}_{1}$ versus $\mathcal{D}_{2}$, then we would have stronger evidence that LLMs might not be distracted by the presence of plausibly memorized causal relations. Table 2 summarizes the results across the two benchmarks. We see that the previously top-performing GPT models are all affected by distracting causal relations when considering effects in the bivariate DAG, with GPT-4-Turbo’s accuracy decreasing to 55%. Interestingly, these models fare better in the mediation DAG setting. However, these findings point to the significant impacts that distracting aspects of prompts can have on LLM reasoning, suggesting continued caution if using LLMs for decision-making tasks.
  1. 1.
  2. 2.
Graph Type Bivariate Confounding Mediation
Intervened Variable B B C B C
GPT-3.5 IE = 0 $0.93\pm 0.06$ $0.6\pm 0.13$ $0.47\pm 0.13$ $0.33\pm 0.12$ $0.67\pm 0.12$
IE = 1 $0.8\pm 0.1$ $0.67\pm 0.12$ $0.2\pm 0.1$ $0.37\pm 0.12$ $0.67\pm 0.12$
GPT-4 IE = 0 $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$ $0.87\pm 0.09$ $1.0\pm 0.0$
IE = 1 $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$ $0.8\pm 0.1$ $0.93\pm 0.06$
IE = 1 $0.93\pm 0.06$ $0.93\pm 0.06$ $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$
Table 3: IE prediction performance across sub-cases to isolate the effect of relation retrieval on Random $(\mathcal{D}_{1})$ benchmark. LLMs do not significantly rely on shortcuts related to relation retrieval from the input prompt. Since intervening on variable A never changes the causal graph in any scenario, we do consider them for this analysis and leave them blank.
Note that drop in performance on group ($IE=1$) as compared to the group ($IE=0$) can indicate reliance on shortcuts based on relation retrieval. We consider the average performance of the LLMs on the effects in each group, selecting only those LLMs which achieve an accuracy $\geq 0.95$ on relation retrieval (which we report in Appendix Table 9). We focus on the random benchmark ($\mathcal{D}_{1}$) to exclude any impacts due to memorized variable names. Table 3 summarizes this study. We find that the general trend does not show strong reliance on shortcuts across LLMs; only GPT-3.5 has a significant drop in its relative performance on group ($IE=1$) vs group ($IE=0$) for the confounding DAG case.
Graph Type Bivariate Confounding Mediation
Intervened Variable A B A B C A B C
Table 4: IE prediction accuracy on the Random $(\mathcal{D}_{1})$ benchmark for the substitution task. The performance of GPT-3.5 & GPT-4 is worse under substitution compared to the non-substitution case (Table 1), while GPT-4-turbo does not show significant change.
RQ4: Are LLMs robust to descriptions of interventions in-context?  Consider the mapping function $M_{I}(\cdot)$ that verbalizes an intervention in Figure 2. We ask whether LLMs can achieve the same performance on the intervention effects in the random benchmark if we varied the mapping function $M_{I}(\cdot)$ to instead “teach” an LLM (Patel et al., 2023) about a new graphical operation that behaves identically to an intervention. We randomly generate strings to instantiate this operation. Figure 3 in the appendix illustrates how we generate prompts for this task which we refer to as the substitution task. Table 4 summarizes IE prediction accuracy for the substitution task on the Random benchmark $\mathcal{D}_{1}$. Contrasting these results against Table 1, we see that for GPT-4 and GPT-3.5, the performance generally suffers. However, GPT-4-turbo appears to be robust to changes in the way interventions are described.
5 Related Work
This paper relates most closely to recent papers that develop benchmarks to evaluate LLMs on various causal reasoning tasks. Kıcıman et al. (2023) introduced multiple causal reasoning benchmarks for LLMs, including evaluating the ability of LLMs to recover the bivariate causal DAGs introduced in the Tübingen pairs dataset (Mooij et al., 2016). Kıcıman et al. (2023) found that GPT models recovered known causal relationships with up to 96% accuracy when experimenting with various prompting strategies such as including system prompts. Building on these results, several recent papers use LLMs to augment the discovery of causal DAGs using natural language descriptions of variables (Abdulaal et al., 2024; Castelnovo et al., 2024; Jiralerspong et al., 2024; Takayama et al., 2024; Long et al., 2023b; a; Ban et al., 2023; Vashishtha et al., 2023). However, evaluating LLMs on their ability to retrieve causal knowledge about known variables constitutes commonsense causal reasoning. In contrast, this paper contributes to work that evaluates abstract causal reasoning (Binz & Schulz, 2023; Jin et al., 2024; 2023; Cai et al., 2023; Liu et al., 2024), assessing the ability of LLMs to use axioms of causality to solve tasks involving general or even new variables.
In the vein of causal reasoning, Jin et al. (2023) studied whether LLMs could correctly infer some causal relationships based on conditional independence statements, comparing LLM predictions to those made by an oracle causal discovery algorithm for observational data, where all causal relations cannot be resolved. In contrast to observational causal discovery, this paper focuses on reasoning about interventions. Also focusing on interventions, Jin et al. (2024) introduced CLadder, a comprehensive benchmark that includes the estimation of causal effects from quantitative data. Causal effect estimation is a complex task that requires solving multiple sub-tasks such as: (i) parsing the prompt to extract a causal DAG, (ii) inferring a function that estimates the effect given the DAG, and (iii) applying that function to the given quantitative data. Concurrently, Cai et al. (2023) introduced a task that asks LLMs to output only causal relationships given a tabular dataset that includes variable names. They focus on disentangling the impacts that prior knowledge (e.g., variables names) and quantitative data have on LLM performance. The empirical study we conduct to assess whether LLMs are sensitive to the presence of plausibly memorized causal relations is similar to experiments conducted by Cai et al. (2023). In contrast to these benchmarks, intervention effects target a narrower question than the general estimation of causal effects, since intervention effects involve binary classification only (i.e., the absence/presence of causal relations in DAGs). We argue that the evaluations we design better isolate causal reasoning from sub-tasks like drawing statistical inferences from quantitative data provided in-context, which both CLadder and the work of Cai et al. (2023) require.
In focusing on intervention effects, we build on the work of Binz & Schulz (2023), who were motivated by prior work in psychology (Waldmann & Hagmayer, 2005) that shows humans weight collected observational evidence and experimental evidence differently when drawing causal conclusions. Binz & Schulz (2023) adapted this psychology study for LLMs, creating prompts that describe observational and post-interventional findings to LLMs to see if they update their beliefs about a system after interventions. They found that GPT-3, unlike human subjects, fared poorly at understanding the implications of interventions. Motivated by their focus on intervention-based reasoning, we significantly expand on the evaluations designed by Binz & Schulz (2023), systematically generating intervention effects with varying degrees of difficulty to further explore the effects of plausible memorization and shortcuts like relation retrieval.
Finally, we note that recent work (Zečević et al., 2023) has called into question whether LLMs are causal parrots, mimicking behavior seen during training without being able to generalize the use of causal logic. Kıcıman et al. (2023); Cai et al. (2023); Jin et al. (2024) as well as the benchmarks designed in this paper aim to stress-test LLMs more extensively by varying prompt generation to assess different shortcuts that LLMs could have learned.
@@ 6 Discussion and Limitations
The goal of this paper is to introduce a causal reasoning benchmark that stress-tests the ability of LLMs to accurately predict how knowledge should be updated after interventions are performed, without conflating other aspects of reasoning such as statistical inference on quantitative data. The research questions that we investigated point to both some optimism and caution. While on one hand, in some scenarios, GPT-4 appears to accurately predict how interventions modify given causal relations, on the other hand, its performance can be negatively affected when prompts describe causal knowledge that it has plausibly memorized from training. Overall, these findings point to the continued importance of designing benchmarks and studies that evaluate varied aspects of abstract causal reasoning in LLMs, especially if practitioners wish to use LLMs to generate candidate decisions.
@@
Finally, keeping in mind Goodhart’s law 444en.wikipedia.org/wiki/Goodhart’s_law, we stress that evaluations like ours and those of others do not serve as metrics for LLMs to beat before being deployed in high-stakes situations. We discourage gaming these benchmarks, and instead intend for this study to, like psychology studies, shed light on LLM behavior.
@@ Acknowledgements
We would like to thank Tejas Vaidhya for contributing to experiments during the initial phase of the project. This work is supported by CIFAR and enabled in part by resources from Digital Research Alliance of Canada (https://alliancecan.ca), Mila (https://mila.quebec), and NVIDIA.
@@ Appendix A Experiment Setup Details
A.1 Dataset Generation
Random ($\mathcal{D}_{1}$).
For causal graphs in this benchmark, the graph nodes are named as some arbitrary string of characters (E.g. ’xyz’) that are independently generated for each graph.
Tübingen ($\mathcal{D}_{2}$).
@@ A.1 Dataset Generation
  • Bivariate. We randomly sample cause-effect pairs from the TP dataset and each cause-effect pair is used to generate a graph.
@@
  • Confounding. We randomly sample two different cause effect pairs such that both have the common cause.
    Example - Altitude (A) - Temperature (B) and Altitude (A) - Precipitation (C) are cause-effect pairs in the TP dataset, and the input graph states: Altitude causes temperature ; Altitude causes precipitation.
  • Mediation. We randomly sample two different cause effect pairs such that the effect of one pair is the cause in the other pair.
    Example - Altitude (A) - Temperature (B) and Temperature (B) - Ozone concentration (C) are cause-efffect pairs in the TP dataset, and the input graph states: Altitude causes temperature ; Temperature causes Ozone concentration.
Anti-commonsense ($\mathcal{D}_{3}$).
Similar to the Tübingen ($\mathcal{D}_{2}$) case, we select the graph nodes from the TP dataset. However, we change the input causal relations between them in an adversarial manner as follows:
  • Bivariate. We randomly sample a cause-effect pairs from the TP dataset and the provide the reversed cause-effect relationship in the input graph.
    Example - Altitude (A) - Temperature (B) is a cause-effect pair in the TP dataset, and the input graph states: Temperature causes altitude.
  • Confounding. For variable $B$ and $C$, we first randomly sample cause effect pairs and then for sampled pair we randomly select a variable (for $A$) from the TP dataset that it is different from the variables in the corresponding cause-effect pair.
@@
A.2 Details regarding LLMs
Querying LLMs.
To query the GPT models, we used the OpenAI API  [^5] https://openai.com/blog/openai-api through the Langchain interface  [^6] https://python.langchain.com/docs/expression_language/interface. Meanwhile, Hugging Face’s Transformers API [^7] https://huggingface.co/docs/transformers/en/index was used to load and query LLaMA-2.
Regarding compute resources, since the GPT models are hosted remotely by OpenAI, we were able to make API calls locally with minimal CPU usage. Conversely, for LLaMA-2, we first had to load the model onto a cluster equipped with a GPU (RTX 8000/32 GB RAM) before submitting our queries.
LLM Output.
@@ A.2 Details regarding LLMs
A.3 Substitution Task
Figure 3: Prompt design for intervention reasoning in the substitution task.
For this task, we first describe the concept of interventions to LLM and refer to them as some arbitrary string of characters (E.g. ’xyz’). We then follow the same prompt template (Figure 2) to test the intervention effect performance, except we never use the word intervention again and replace it with the chosen substitution word.
A.4 Ground Truth Intervention Effects
This should help the reader understand clearly what queries we considered for the different research questions in our analysis. For the case of RQ3, we partition the queries into two sets based on the value of ground truth IE being either 0 or 1. While for the case of RQ2, we considered both the queries for the Bivariate case (intervention on B); while for the Mediator case (intervention on either B or C) we only considered the query ($A\rightarrow C$).
@@ B.1 Reproducing results on the Tübingen dataset
We reproduced the results of Kıcıman et al. (2023) on the Tübingen dataset as shown in Table 6. Since their task did not involve the LLM to reason about the effect of interventions, we term this task of inferring relations from the input graph in the prompt as relation retrieval.
@@
B.2 RQ1: How accurate are LLMs at predicting the effects of interventions?
@@
To understand how effective LLMs are at IE prediction, we focused on the Random ($\mathcal{D}_{1}$) benchmark in the main paper (Table 1), as we want to remove any distracting effect of semantically meaningful entities as graph nodes. We now present the same results on the Tübingen $(\mathcal{D}_{2})$ and the Anti-commonsense $(\mathcal{D}_{3})$ benchmark as well in Table 7.
@@
We find that the results in both the cases are very similar to the case with the Random ($\mathcal{D}_{1}$) benchmark; GPT-4-turbo performs the best and GPT models outperform LLaMA-2 by a big margin.

### v2

@@ 5 Empirical Analysis
RQ2: To what extent is LLM performance affected by possibly memorized causal relations?  While we might be tempted to conclude that GPT-4 reliably predicts changes to models after interventions are performed, we consider spurious factors that can affect model performance. In particular, Kiciman et al. (2024) found that GPT models reliably retrieved information about TP causal relations, suggesting that these relations could have been included in the training data for LLMs. (We reproduce their findings in Appendix Table 4.) This leads to a worrying possibility: after interventions, could LLMs fail to update their beliefs about causal relations that they have potentially memorized? To study this, we consider only the subset of causal relations that appear in TP and the interventions that sever these relations. If an LLM achieves good performance in general by having memorized some known causal relations, then it would achieve worse performance on Tübingen for the selected IEs, which sever known causal relations and go against the LLM’s learned biases, compared its performance on Random or Random Tübingen , where post-interventional causal relations do not directly contradict known facts. In Table 6, we specify which causal relations appear in TP for each causal DAG we study, and which interventions modify these relations. Table 2 summarizes the results. Bolded figures indicate that the performance on Tübingen drops significantly (with $\alpha=0.05$) compared to performance on either of the other benchmarks. Interestingly, GPT do not show evidence of relying purely on memorized causal relations, since performance is overall comparable across the benchmarks. Further, the models seem to struggle with the Random Tübingen benchmark more than they do with the other benchmarks, leading to a question about whether they are in fact sensitive to variations on the TP dataset.
Table 2: IE prediction accuracy considering only causal relationships that the LLM could potentially memorize. For causal relations that appear in TP, and for interventions that modify these relations, worse performance on the Tübingen benchmark compared to performance on Random Char or Random Tübingen provides evidence that LLMs might be relying on memorized facts to achieve high accuracy on IEs. Bolded figures indicate performances that are significantly ($\alpha=0.05$) worse than the corresponding performance on Tübingen .
Model Graph Intervened Variable Random Char Tübingen Random Tübingen
GPT-3.5 Bivariate B $0.8\pm 0.1$ $0.8\pm 0.1$ $0.53\pm 0.13$
Confounding C $0.47\pm 0.13$ $0.4\pm 0.13$ $0.47\pm 0.13$
Mediation B $0.4\pm 0.13$ $0.4\pm 0.13$ $0.4\pm 0.13$
C $0.67\pm 0.12$ $0.67\pm 0.12$ $0.6\pm 0.13$
GPT-4 Bivariate B $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$
Confounding C $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$
Mediation B $0.8\pm 0.1$ $\bf{0.73\pm 0.11}$ $0.93\pm 0.06$
C $0.93\pm 0.06$ $1.0\pm 0.0$ $0.93\pm 0.06$
GPT-4-turbo Bivariate B $0.93\pm 0.06$ $0.8\pm 0.1$ $0.73\pm 0.11$
Confounding C $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$
Mediation B $1.0\pm 0.0$ $1.0\pm 0.0$ $0.93\pm 0.06$
C $1.0\pm 0.0$ $1.0\pm 0.0$ $0.93\pm 0.06$
GPT-4o Bivariate B $1.0\pm 0.0$ $0.93\pm 0.06$ $0.73\pm 0.11$
Confounding C $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$
Mediation B $1.0\pm 0.0$ $1.0\pm 0.0$ $0.87\pm 0.09$
C $0.93\pm 0.06$ $0.87\pm 0.09$ $0.8\pm 0.1$
@@ 6 Discussion and Limitations
The goal of this paper is to introduce a causal reasoning benchmark that stress-tests the ability of LLMs to accurately predict how knowledge should be updated after interventions are performed, without conflating other aspects of reasoning such as statistical inference on quantitative data. Our findings are optimistic, but we believe that nevertheless, they underscore the continued need for benchmarks and studies that evaluate varied aspects of abstract causal reasoning in LLMs, especially if practitioners wish to use LLMs to generate candidate decisions.
@@ Acknowledgements
We would like to thank Tejas Vaidhya for contributing to experiments during the initial phase of the project. DS acknowledges support from NSERC Discovery Grant RGPIN-2023-04869, and a Canada-CIFAR AI Chair.
@@ Appendix A Experiment Setup Details
A.1 Causal DAGs
We consider 3 basic causal DAGs: Bivariate, Confounding and Mediation as shown in Figure 2
A.2 Dataset Generation
Random Char (R).
In the causal graphs used in this benchmark, the nodes are labeled with arbitrary English letters (e.g. ’z’, ’b’) that are independently generated for each graph.
Tübingen (T).
@@ A.2 Dataset Generation
  • Bivariate. We sample cause-effect pairs from the TP dataset and each cause-effect pair is used to generate a graph.
@@
  • Confounding. For variable $B$ and $C$, we first sample cause-effect pairs and then for sampled pair we randomly select a variable (for $A$) from the TP dataset that it is different from the variables in the corresponding cause-effect pair.
@@
The rationale behind defining the Tübingen dataset in this manner is elaborated in Table 6.
Random Tübingen (RT).
Similar to the Tübingen case, we select the graph nodes from the set of variables present in the TP dataset. However, none of the causal relations defined in the graphs are present in the TP dataset. Instead the causal relations defined are between randomly selected variables, as follows:
  • Bivariate. We randomly sample two unrelated variables from the TP dataset and define a cause-effect relationship in the input graph.
    Example - Age ($A$) - Temperature ($B$) are variables without a cause-effect relation in the TP dataset, and the input graph states: Age causes Temperature.
  • Confounding. We randomly sample 3 variables from the TP dataset that no two variables selected have a causal relation in the TP dataset and we define the graph.
    Example - Altitude (A) - Horsepower (B) - Cement (C) are randomly selected variables from the TP dataset. No two variables have a causal relationship in the TP dataset. The relationships in the input graph is as follows: Altitude causes Horsepower ; Altitude causes Cement.
  • Mediation. We select the 3 variables exactly as we did for the confounding case and define the graph.
    Example - Altitude (A) - Horsepower (B) - Cement (C) are randomly selected variables from the TP dataset. The relationsihips in the input graph is as follows: Altitude causes Horsepower ; Horspower causes Cement.
A.3 Details regarding LLMs
A.3.1 Querying LLMs.
To query the GPT models, we used the OpenAI API  [^1] https://openai.com/blog/openai-api through the Langchain interface  [^2] https://python.langchain.com/docs/expression_language/interface. Meanwhile, VLLM library 333hhttps://docs.vllm.ai/en/latest/ was used for fast inference from LLaMA models.
Regarding compute resources, since the GPT models are hosted remotely by OpenAI, we were able to make API calls locally with minimal CPU usage. Conversely, for LLaMA models, we first had to load the model onto a cluster equipped with a GPU (A100/80 GB RAM) before submitting our queries.
A.3.2 LLM Output.
@@ A.3 Details regarding LLMs
  • LLaMA-3: llama-3.1-8b
A.4 Substitution Task
Figure 3: Prompt design for intervention reasoning in the substitution task. Instead of using the word intervention directly in the prompt, we illustrate the concept of interventions and define it with a random word, for instance, operation ’xyz’ in the prompt template above.
For this task, we first describe the concept of interventions to LLM and refer to them as some arbitrary string of characters (E.g. ’xyz’). We then follow the same prompt template (Figure 1) to test the intervention effect performance, except we never use the word intervention again and replace it with the chosen substitution word. The prompt verbalization for this task is illustrated in Figure 3.
A.5 Ground Truth Intervention Effects
@@ C $A\to B$ 0
This should help the reader understand clearly what queries we considered for the different research questions in our analysis.
@@ Appendix B Additional Results
We now go over the empirical studies in detail and provide the results for the remaining research questions.
@@ B.1 Reproducing results on the Tübingen dataset
We reproduced the results of Kiciman et al. (2024) on the Tübingen dataset as shown in Table 4. Since their task did not involve the LLM to reason about the effect of interventions, we term this task of inferring relations from the input graph in the prompt as relation retrieval.
@@
Table 5: Relation Retrieval accuracy of models on Random benchmark. The GPT models show good performance but LLaMA-2 performs poorly. Hence, as per our criteria we drop the LLaMA-2 model for analysis in RQ3.
Model Bivariate Confounding Mediation
GPT-3.5 $1.0\pm 0.0$ $1.0\pm 0.0$ $0.98\pm 0.02$
GPT-4 $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$
GPT-4-turbo $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$
GPT-4o $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$
LLaMA-2 $0.73\pm 0.11$ $0.59\pm 0.13$ $0.84\pm 0.09$
LLaMA-3 $1.0\pm 0.0$ $1.0\pm 0.0$ $0.94\pm 0.06$
B.2 Research Questions
@@
To understand how effective LLMs are at IE prediction, we focused on the Random (R) benchmark in the main paper (Table 1), as we wanted to remove any distracting effect of semantically meaningful entities as graph nodes. We now present the same results on the Tübingen (T) and Random Tübingen (RT) benchmarks as well in Table 6.
@@
We find that the results in both the cases are very similar to the case with the Random benchmark; GPT-4-turbo performs the best and GPT models outperform LLaMA models by a big margin.
Table 6: IE prediction accuracy on all three benchmarks. GPT-4 variants are the best performing models, while LLaMA models struggle with interventional reasoning.
Graph Type Benchmark Bivariate Confounding Mediation
Intervened Variable A B A B C A B C
GPT-3.5 R $0.83\pm 0.08$ $0.87\pm 0.06$ $0.8\pm 0.09$ $0.69\pm 0.12$ $0.36\pm 0.09$ $0.58\pm 0.11$ $0.36\pm 0.12$ $0.67\pm 0.12$
T $0.87\pm 0.06$ $0.83\pm 0.06$ $0.82\pm 0.05$ $0.67\pm 0.11$ $0.31\pm 0.07$ $0.64\pm 0.08$ $0.42\pm 0.1$ $0.67\pm 0.12$
RT $0.83\pm 0.1$ $0.7\pm 0.12$ $0.87\pm 0.09$ $0.67\pm 0.12$ $0.47\pm 0.13$ $0.64\pm 0.12$ $0.33\pm 0.12$ $0.63\pm 0.12$
GPT-4 R $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$ $0.78\pm 0.09$ $0.82\pm 0.08$ $0.96\pm 0.03$
T $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$ $0.84\pm 0.07$ $1.0\pm 0.0$
RT $1.0\pm 0.0$ $1.0\pm 0.0$ $0.93\pm 0.06$ $0.87\pm 0.09$ $1.0\pm 0.0$ $0.98\pm 0.04$ $0.97\pm 0.05$ $0.97\pm 0.05$
GPT-4-turbo R $1.0\pm 0.0$ $0.97\pm 0.03$ $0.96\pm 0.03$ $0.93\pm 0.05$ $1.0\pm 0.0$ $0.98\pm 0.02$ $1.0\pm 0.0$ $1.0\pm 0.0$
T $0.93\pm 0.04$ $0.9\pm 0.05$ $0.62\pm 0.09$ $0.78\pm 0.08$ $0.96\pm 0.03$ $0.69\pm 0.09$ $0.96\pm 0.04$ $0.96\pm 0.04$
RT $0.8\pm 0.1$ $0.8\pm 0.1$ $0.73\pm 0.11$ $0.6\pm 0.13$ $1.0\pm 0.0$ $0.8\pm 0.1$ $0.97\pm 0.05$ $0.97\pm 0.05$
GPT-4o R $1.0\pm 0.0$ $1.0\pm 0.0$ $0.98\pm 0.02$ $0.91\pm 0.04$ $1.0\pm 0.0$ $0.93\pm 0.05$ $1.0\pm 0.0$ $0.87\pm 0.06$
T $0.87\pm 0.06$ $0.97\pm 0.03$ $0.89\pm 0.04$ $0.93\pm 0.03$ $1.0\pm 0.0$ $0.84\pm 0.05$ $0.96\pm 0.04$ $0.87\pm 0.06$
RT $0.87\pm 0.09$ $0.83\pm 0.1$ $0.9\pm 0.08$ $0.8\pm 0.1$ $1.0\pm 0.0$ $0.93\pm 0.06$ $0.93\pm 0.06$ $0.87\pm 0.09$
LLaMA-2 R $0.5\pm 0.12$ $0.4\pm 0.12$ $0.56\pm 0.11$ $0.53\pm 0.11$ $0.16\pm 0.06$ $0.69\pm 0.09$ $0.56\pm 0.12$ $0.64\pm 0.12$
T $0.4\pm 0.12$ $0.2\pm 0.08$ $0.42\pm 0.11$ $0.4\pm 0.11$ $0.09\pm 0.04$ $0.47\pm 0.12$ $0.31\pm 0.12$ $0.44\pm 0.1$
RT $0.43\pm 0.13$ $0.23\pm 0.11$ $0.47\pm 0.13$ $0.47\pm 0.13$ $0.2\pm 0.1$ $0.62\pm 0.13$ $0.4\pm 0.13$ $0.7\pm 0.12$
LLaMA-3 R $0.8\pm 0.1$ $0.83\pm 0.1$ $0.47\pm 0.13$ $0.6\pm 0.13$ $0.87\pm 0.09$ $0.38\pm 0.13$ $0.5\pm 0.13$ $0.47\pm 0.13$
T $0.73\pm 0.11$ $0.87\pm 0.09$ $0.53\pm 0.13$ $0.53\pm 0.13$ $1.0\pm 0.0$ $0.38\pm 0.13$ $0.6\pm 0.13$ $0.27\pm 0.11$
RT $0.67\pm 0.12$ $0.87\pm 0.09$ $0.47\pm 0.13$ $0.4\pm 0.13$ $1.0\pm 0.0$ $0.4\pm 0.13$ $0.53\pm 0.13$ $0.43\pm 0.13$
RQ2: To what extent is LLM performance affected by memorized causal relations?  In RQ1, we computed the IE performace by considering all the causal relationship queries for each of the interventions. In the case of RQ2, we only consider specific causal relations that could be potentially memorized by the LLMs and interventions that sever these relations. For:
  1. 1.
    Bivariate DAG, we consider $A\rightarrow B$ with the intervention on $B$.
  2. 2.
    Confounding DAG, we consider $B\rightarrow C$ with the intervention on $C$.
  3. 3.
    Mediation DAG, we consider $A\rightarrow C$ with an intervention on $B$ and $C$ separately.
The Tübingen benchmark (Section A.2) is defined in such a way that, all of the causal relations under consideration above are cause-effect pairs present in the TP dataset that could be potentially memorized by the LLM.
We provide results for role of memorization in LLaMA models in Table 7. Given their overall poor IE performance in the random benchmark, it is difficult to draw conclusions about the impact of memorization but they have similar performance across the benchmarks.
Table 7: IE prediction performance of LLaMA models on specific scenarios for all the benchmarks to understand the role memorization. LLaMA models perform poorly but demonstrate relatively similar performance across the three benchmarks.
Model Graph Intervened Variable Random Tübingen Random Tübingen
LlaMA-2 Bivariate B $0.4\pm 0.13$ $0.4\pm 0.13$ $0.27\pm 0.11$
Confounding C $0.0\pm 0$ 0.0 $0.07\pm 0.06$ $0.33\pm 0.12$
Mediation B $0.33\pm 0.12$ $0.33\pm 0.12$ $0.4\pm 0.13$
C $0.53\pm 0.13$ $0.53\pm 0.13$ $0.6\pm 0.13$
LlaMA-3 Bivariate B $0.67\pm 0.12$ $0.73\pm 0.11$ $0.8\pm 0.1$
Confounding C $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$
Mediation B $0.67\pm 0.12$ $0.53\pm 0.13$ $0.67\pm 0.12$
C $0.53\pm 0.13$ $0.33\pm 0.12$ $0.4\pm 0.13$
Table 8: IE prediction performance across sub-cases to isolate the effect of relation retrieval on Random benchmark. LLMs do not significantly rely on shortcuts related to relation retrieval from the input prompt. We see a significant difference only for the confounding DAG case in GPT-3.5 and for bivariate DAG case in LLaMA-3. Since intervening on variable A never changes the causal graph in any scenario, we don’t consider them for this analysis. Bolded figures indicate performances that are significantly ($\alpha=0.05$) different worse in IE = 1 compared to IE = 0.
Graph Type Bivariate Confounding Mediation
Intervened Variable B B C B C
GPT-3.5 IE = 0 $0.93\pm 0.06$ $0.6\pm 0.13$ $\bf{0.47\pm 0.13}$ $0.33\pm 0.12$ $0.67\pm 0.12$
IE = 1 $0.8\pm 0.1$ $0.67\pm 0.12$ $\mathbf{0.2\pm 0.1}$ $0.37\pm 0.12$ $0.67\pm 0.12$
GPT-4 IE = 0 $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$ $0.87\pm 0.09$ $1.0\pm 0.0$
IE = 1 $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$ $0.8\pm 0.1$ $0.93\pm 0.06$
IE = 1 $0.93\pm 0.06$ $0.93\pm 0.06$ $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$
GPT-4o IE = 0 $1.0\pm 0.0$ $0.87\pm 0.09$ $1.0\pm 0.0$ $1.0\pm 0.0$ $0.87\pm 0.09$
IE = 1 $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$ $0.87\pm 0.09$
LlaMA-2 IE = 0 $0.4\pm 0.13$ $0.5\pm 0.13$ $0.0\pm 0$ 0.0 $0.33\pm 0.12$ $0.67\pm 0.12$
IE = 1 $0.4\pm 0.13$ $0.53\pm 0.13$ $0.13\pm 0.09$ $0.33\pm 0.12$ $0.6\pm 0.13$
LlaMA-3 IE = 0 $\bf{1.0\pm 0.0}$ $0.4\pm 0.13$ $1.0\pm 0.0$ $0.6\pm 0.13$ $0.27\pm 0.11$
IE = 1 $\bf{0.67\pm 0.12}$ $0.6\pm 0.13$ $0.87\pm 0.09$ $0.5\pm 0.13$ $0.47\pm 0.13$
  1. 1.
  2. 2.
Table 3 clearly categorizes the causal relations under intervention into these two groups.
Note that drop in performance on group ($IE=1$) as compared to the group ($IE=0$) can indicate reliance on shortcuts based on relation retrieval. We consider the average performance of the LLMs on the effects in each group, selecting only those LLMs which achieve an accuracy $\geq 0.95$ on relation retrieval (which we report in Appendix Table 5). We focus on the Random benchmark to exclude any impacts due to memorized variable names. Table 8 summarizes this study. We find that the general trend does not show strong reliance on shortcuts across LLMs; GPT-3.5 only has a significant drop in its relative performance on group ($IE=1$) vs group ($IE=0$) for the confounding DAG cases. And, significant difference for bivariate DAG cases in LLaMA-3. Since, LLaMA-2 doesn’t satisfy our constraint of high relation retrieval accuracy, we don’t consider it for further analysis.
RQ4: Are LLMs robust to descriptions of interventions in-context?  Consider the verbalization of an intervention in Figure 1. It mentions the $\mathrm{do}(\cdot)$ operator and refers to a “perfect intervention” to prompt an LLM to rely on facts about causal reasoning that could have appeared in the training dataset. We ask whether LLMs can achieve the same performance on the intervention effects in the random benchmark if we varied the verbalization to instead “teach” an LLM (Patel et al., 2023) about a new graphical operation that behaves identically to an intervention. We randomly generate strings to instantiate this operation. Figure 3 in the appendix illustrates how we generate prompts for this task which we refer to as the substitution task. Table 9 summarizes IE prediction accuracy for the substitution task on the Random benchmark. Contrasting these results against Table 1, we see that for the GPT-3.5, GPT-4o, and LLaMA models, performance generally suffers. However, GPT-4-turbo appears to be robust to changes in the way interventions are described.
Table 9: IE prediction accuracy on the Random benchmark for the substitution task. Except for GPT-4 and GPT-4-turbo, rest of the models perform worse under substitution compared to the non-substitution case (Table 1).